The probability that the sum is 5 when you roll two dice is 1/9; the probability...

  1. The probability that the sum is 5 when you roll two dice is 1/9; the probability that the sum is 6 when you roll two dice is 5/36. Suppose you play a game where you win if the sum is 5 or 6. What is the probability that you win?

    A. 1/2 B. 1/4 C. 2/3 D. 6/45 E. 6/36

  2. Suppose a sorority is selling 650 raffle tickets to raise money for a local charity. You are going to be the first person to purchase any of the tickets. What is the smallest number of tickets you need to purchase to ensure that your probability of winning is greater thananyone else’s probability of winning?

    A. 650 B. 424 C. 326 D. 325 E. 2

  1. If individual outcomes are uncertain but there is a regular distribution of outcomes in a large number of repetitions, we call the phenomenon:

    A. Random B. Probable C. Predictable D. Unlikely
Answer #1

1) probability that you win =(1/9)+(5/36)=9/36 =1/4

2)

smallest number of tickets you need to purchase to ensure that your probability of winning is greater thananyone else’s probability of winning =(650/2)+1=326

option C

3) A. Random